OpenStudy (anonymous):

2sinAsinB=cos(A-B)-cos(A+B)

OpenStudy (mertsj):

\[\cos (A-B)-\cos (A+B)=\] \[\cos A \cos B+\sin A \sin B-(\cos A \cos B-\sin A \sin B)\]

OpenStudy (mertsj):

You can finish it

OpenStudy (anonymous):

2sinAsinB ?

OpenStudy (mertsj):

What were you supposed to show that it is equal to?